# A Systematic Survey on Large Language Models for Evolutionary Optimization: From Modeling to Solving

### v2

@@ IV-C Challenges
With the growing application of LLMs in optimization, increasing efforts have been made to bridge natural language descriptions of optimization problems with formal mathematical models. Frameworks such as Chain-of-Experts [70] and ORLM [27] demonstrate that LLMs can not only complete modeling tasks efficiently but also uncover implicit constraints that may be overlooked by human experts, thus introducing a new paradigm for optimization modeling. Nevertheless, existing research still faces several key challenges:
  • Prompt-based methods are widely adopted due to their ease of deployment and the absence of large annotated datasets. However, their reliance on closed-source LLM APIs raises serious concerns regarding data security and privacy, especially in industrial settings involving sensitive information. Moreover, even advanced models such as GPT-4 exhibit limitations in mathematical reasoning, leading to issues of accuracy and robustness on fundamental modeling tasks [27]. Hence, enhancing the intrinsic reasoning capability of LLMs remains a primary challenge for this paradigm.
  • Learning-based methods such as ReSocratic [88] and OptMATH [91] have investigated synthetic data generation to construct large-scale optimization datasets. Yet, studies suggest that comparable performance can sometimes be achieved using only a fraction of the dataset, pointing to diminishing returns from data scaling [110]. Consequently, an open problem is how to efficiently generate high-quality, representative datasets that maximize learning efficiency. Furthermore, as this line of research is still at an early stage, the number of existing studies remains limited, and broader exploration is needed to fully exploit the potential of learning-based modeling.
@@ V LLMs for Optimization Solving
Optimization solving is a core stage of the optimization workflow, and decades of research have produced a wide range of high-quality algorithms. In recent years, machine learning methods have been increasingly integrated into optimization to improve performance, and the advent of LLMs has further propelled this research into a new phase. In this section, we categorize existing work into three paradigms: LLMs as optimizers, low-level LLM-assisted optimization algorithms, and high-level LLM-assisted optimization algorithms. Section V-A introduces LLMs as optimizers, which directly employ LLMs to solve problems without relying on traditional frameworks (see Fig. 4). Sections V-B and V-C discuss how LLMs can enhance optimization algorithms at the component level and at the orchestration/design level, respectively (illustrated in Figs. 5 and 6). Section V-D then outlines the key challenges in this area. Representative studies are summarized in Table III.
Method Venue Type Technical Summary LLMs as Optimizers OPRO [28] ICLR, 2023 Prompt-based Iteratively refine through optimization trajectories and problem descriptions. toLLM [111] KDD, 2023 Prompt-based Design four canonical tasks to evaluate the performance boundaries of LLMs. EvoLLM [112] GECCO, 2024 Prompt-based Replace traditional optimization trajectories with candidate solution quality ranking. MLLMO [113] MCII, 2024 Prompt-based Utilize MLLMs to jointly process problem descriptions and map visualizations for CVRPs. POM [114] NeurIPS, 2024 Learning-based Pre-train a general-purpose, zero-shot black-box optimization foundation model. OPTO [115] NeurIPS, 2024 Prompt-based Replace traditional optimization trajectories with rich, structured execution traces. VRMA [116] arXiv, 2024 Prompt-based Utilize MLLMs to process 2D planar point distribution maps as input. ROPRO [117] ACL, 2024 Prompt-based Discuss the model dependency of OPRO and identify its limitations on small-scale models. BBOLLM [29] arXiv, 2024 Prompt-based Conduct evaluation of LLMs on both discrete and continuous black-box optimization problems. LLMS [118] arXiv, 2024 Learning-based Fine-tune LLMs with instruction-solution pairs for scheduling problems. ECLIPSE [119] NAACL, 2024 Prompt-based Apply iterative optimization to jailbreaking attack strategies. LLOME [120] arXiv, 2024 Learning-based Integrate preference learning to train LLMs for satisfying complex biophysical constraints. LLMDSM [121] arXiv, 2024 Prompt-based Apply iterative optimization to design structure matrix sequencing. LMCO [122] WCL, 2024 Prompt-based Apply iterative optimization to wireless network design. MGSCO [123] arXiv, 2025 Prompt-based Process visual representations of abstract graph structures using MLLMs. ORFS [124] arXiv, 2025 Prompt-based Apply iterative optimization to automated parameter tuning in chip design. Low-level LLMs for Optmization Algorithms LMX [125] TELO, 2023 Operators Leverage LLMs as intelligent operators for textual genome crossover and recombination. GPT-NAS [126] arXiv, 2023 Initialization Utilize LLMs for NAS initialization with prior knowledge. LMEA [127] CEC, 2023 Operators Employ LLMs as crossover, mutation, and selection operators to guide EAs. LLM-PP [128] arXiv, 2023 Initialization Utilize LLMs as performance predictors to assist initialization processes. LMOEA [129] EMO, 2023 Operators Empower MOEA/D with LLMs through zero-shot prompting as search operators in multi-objective optimization. OFPLLM [130] AI, 2023 Initialization Assist non-expert users in initializing financial plans. LEO [131] Neucom, 2024 Operators Utilize LLMs to guide individuals from dual pools for exploration and exploitation. LLM-GA [132] BiB, 2024 Initialization Initialize genetic algorithms with LLMs to generate high-quality mutant pools for enzyme design. GE [133] GECCO, 2024 Operators Enhance the creativity and diversity of the LLM by introducing different role-based prompts for NAS. LLMAES [134] ICIC, 2024 Operators Generate minimal solutions within populations using LLMs to reduce interaction costs. LLMTES [135] arXiv, 2024 Configuration Apply LLM-based feedback loops to sequentially optimize evolution strategies through static tuning. LAEA [136] SWEVO, 2024 Evaluation Transform model-assisted selection tasks into classification and regression problems. LICO [137] arXiv, 2024 Evaluation Apply LLMs as surrogate models in molecular science applications. LLMSM [138] CAI, 2024 Evaluation Develop a collaborative framework using LLMs for model selection and training in engineering optimization. LLMCES [139] GECCO, 2024 Configuration Implement step-size control for evolution strategies using OPRO-like mechanisms. AKFLLM [140] arXiv, 2024 Operators Utilize LLMs to assist mutation and other generative stages in evolutionary multi-task optimization. LLMAMO [141] arXiv, 2024 Operators Invoke LLMs for elite solution generation only upon insufficient population improvement. LTC [142] ESANN, 2024 Configuration Apply LLM-based feedback loops to sequentially optimize CMA-ES through dynamic control. LLMEVO [143] arXiv, 2025 Operators Validate LLM effectiveness in selection, crossover, and mutation phases while noting limitations in initialization. LLM-GE [144] GECCO, 2025 Operators Utilize LLMs as crossover and mutation operators to optimize YOLO architectures for object detection. LAOS [145] GECCO, 2025 Configuration Utilize state features to replace optimization trajectories for guiding LLM-based operator selection. PAIR [146] arXiv, 2025 Operators Focus on utilizing LLMs as selection operators to enhance LMEA. LLMMS [147] arXiv, 2025 Evaluation Utilize LLMs as meta-Surrogates to facilitate cross-task knowledge sharing through token sequence representations. LMPSO [148] arXiv, 2025 Operators Simulate PSO evolutionary processes using LLMs and refine LMEA for specific algorithms. LLM-SAEA [149] arXiv, 2025 Evaluation Dynamically select appropriate proxy models and infill sampling criteria using LLMs. High-level LLMs for Optmization Algorithms LLMGM [150] GECCO, 2023 Single-step Decompose and recombine six existing metaheuristic algorithms to generate novel hybrid algorithms in single rounds. FunSearch [151] Nature, 2023 Iterative Generate novel code snippets via LLMs and guide EAs to search through function space. AEL [152] arXiv, 2023 Iterative Enhance FunSearch by integrating heuristic principles, demonstrating significant improvements on TSP instances. AS-LLM [153] IJCAI, 2023 Selection Automatically extract high-dimensional algorithm features from code or descriptive text to facilitate algorithm selection. EoH [154] ICML, 2024 Iterative Implement heuristic-code coevolution via preset prompt pairs for algorithm crossover and mutation. ReEvo [155] NeurIPS, 2024 Iterative Utilize LLMs as hyper-heuristic algorithms with reflective evolution mechanisms to guide search processes. AutoSAT [156] arXiv, 2024 Iterative Employ multiple heuristic strategies to guide LLMs in algorithm generation. LLaMEA [157] TEVC, 2024 Iterative Utilize LLMs to guide EAs in mutation, selection, and other operations for generating advanced heuristics. L-AutoDA [158] GECCO, 2024 Iterative Apply AEL to generate adversarial attack algorithms in cybersecurity. MOELLM [159] TEVC, 2024 Iterative Introduce robust testing modules and dynamic selection strategies to generate multi-objective optimization algorithms. LLMEPS [160] PPSN, 2024 Iterative Develop a baseline algorithm for automated algorithm design using EoH and ReEvo methodologies. EvolCAF [161] PPSN, 2024 Iterative Apply EoH to evolve and generate cost-aware acquisition functions for Bayesian optimization. TS-EoH [162] arXiv, 2024 Iterative Apply EoH to generate algorithms for edge server task scheduling. MEoH [163] AAAI, 2024 Iterative Extend EoH to multi-objective versions, incorporating other metrics for comprehensive algorithm evaluation. LLaMEA-HPO [164] TELO, 2024 Iterative Integrate hyperparameter optimization into LLaMEA’s iterative cycle to enhance overall efficiency. CMLLM [165] EvoApps, 2024 Iterative Develop dynamic prompt mechanisms to control LLM mutation for enhancing LLaMEA performance. HSEvo [166] AAAI, 2024 Iterative Combine harmony search with GAs and optimize algorithm portfolio quality/diversity through diversity metrics. LLM4AD [167] arXiv, 2024 Iterative Develop an EoH-based platform to accelerate heuristic algorithm design. MCTS-AHD [168] arXiv, 2025 Iterative Employ MCTS to explore heuristic spaces, organizing generated heuristics into tree structures. CEG [169] GECCO, 2025 Iterative Develop methodologies to analyze LLM-generated code and its evolutionary patterns during optimization processes. OPSLAD [170] GECCO, 2025 Iterative Apply LLaMEA to photonic structure optimization in industrial design. BLADE [171] GECCO, 2025 Iterative Establish a standardized and reproducible evaluation framework for LLM-driven algorithm generation methods. AutoHD [172] arXiv, 2025 Iterative Propose an LLM-guided heuristic discovery strategy for complex planning tasks. LAMA [173] TEVC, 2025 Iterative Propose an automatic memetic algorithm enhanced by a heuristic designed with the assistance of LLMs. LAS [174] arXiv, 2025 Iterative Analyze fitness landscapes of LLM-assisted algorithm search processes. LLaMEA-BO [175] arXiv, 2025 Iterative Apply LLaMEA to automatically generate Bayesian optimization algorithms. InstSpecHH [176] arXiv, 2025 Selection Leverage LLMs’ semantic comprehension and reasoning capabilities for context-aware algorithm selection. BSALMD [177] arXiv, 2025 Iterative Conduct behavioral space analysis to examine algorithmic evolution trajectories and behaviors. FLAAD [178] arXiv, 2025 Iterative Fine-tune LLMs with diversity-aware ranked sampling for algorithm generation. EoH-S [179] arXiv, 2025 Iterative Extend EoH to evolve and generate complementary algorithm portfolios.
@@ V-A LLMs as Optimizers
LLMs possess strong in-context learning capabilities, enabling them to solve optimization problems through iterative prompting. A pioneering work in this direction is the OPRO framework proposed by Yang et al. [28], which formulates optimization tasks in natural language and iteratively refines solutions based on historical trajectories. By progressively improving solution quality, OPRO was the first to demonstrate the feasibility of using LLMs as general-purpose optimizers, inspiring subsequent research across diverse domains. Building on this paradigm, Guo et al. [111] benchmarked LLMs on four classical optimization tasks i.e., gradient descent, hill climbing, grid search, and black-box optimization, where the model iteratively generated and evaluated solutions through interactive prompting. To ensure fair and consistent evaluation, they also introduced three task-agnostic metrics designed to reduce sensitivity to test sample variations. Beyond benchmarking, OPRO has been extended to multiple application domains. Jiang et al. [121] applied it to combinatorial optimization for structural matrix ordering; Qiu et al. [122] developed LMCO for wireless network design; Ghose et al. [124] employed it for parameter tuning in open-source hardware workflows; and Jiang et al. [119] adapted the paradigm to jailbreak attack analysis, treating optimizable suffixes and harmfulness scores as part of the optimization trajectory.
@@
In contrast to OPRO’s direct use of solution–feedback trajectories, subsequent research has explored alternative forms of iteration to guide LLMs in optimization tasks. Lange et al. [112] sorted discretized candidate solutions from best to worst and provided the ordered sequence as context to the LLM. This reframes the task from pure solution generation to sequence recognition and reasoning, leveraging the model’s ability to capture ordering patterns. Cheng et al. [115] proposed the OPTO framework, which feeds LLMs with structured execution traces (e.g., intermediate variable values, function calls, and runtime outputs) to simulate backpropagation-like optimization and improve task understanding. Beyond textual iteration, multimodal extensions have further broadened the paradigm by incorporating visual information. Huang et al. [113] addressed vehicle routing problems using both textual descriptions and map images to mimic human spatial cognition. Elhenawy et al. [116] solved the classic traveling salesman problem by combining scatter plot inputs with a multi-agent collaboration mechanism for path search. Zhao et al. [123] extended this idea to graph-structured problems by converting abstract graphs into visual images, enabling multimodal LLMs to capture topological features in a manner more akin to human reasoning. Collectively, these studies demonstrate that extending OPRO with sequence-based, trace-based, or multimodal inputs can substantially enrich the optimization capabilities of LLMs.
Despite the progress of using LLMs directly as optimizers, several inherent drawbacks have been identified. Zhang et al. [117] highlighted the strong model dependency of OPRO, noting that smaller models, with limited reasoning capacity, struggle to exploit iterative histories and generate high-quality solutions. Huang et al. [29] systematically evaluated LLMs on both discrete and continuous black-box optimization problems, confirming their strengths in symbolic and text-based discrete tasks but also exposing their weaknesses in purely numerical settings, such as continuous optimization.
@@
Based on these limitations, researchers have begun to explore alternative avenues for applying LLMs in optimization. One line of work attributes the drawbacks to prompt engineering and the inherent constraints of current models, and thus seeks to improve optimization capability through fine-tuning or training. Abgaryan et al. [118] fine-tuned an LLM to output solutions in a single step, bypassing multi-round iteration by training on instruction–solution pairs. Li et al. [114] introduced POM, a pretrained optimizer model that generalizes across black-box problems in a zero-shot manner. Chen et al. [120] proposed LLOME, a two-level optimization framework that incorporates preference learning to generate biologically valid sequences under complex biophysical constraints. Meanwhile, an increasing body of research suggests that LLMs may be more effective when used to assist rather than replace traditional optimization algorithms. This perspective motivates the next paradigm of low-level LLM-assisted optimization, where LLMs are embedded into existing frameworks to enhance their performance.
@@ V-B Low-level LLMs for Optimization Algorithms
There is a natural complementarity between the collective intelligence of EAs and the knowledge-driven reasoning of LLMs [32]. While EAs excel at heuristic exploration through population-based search, LLMs contribute powerful semantic understanding and generative capabilities that can guide this process, thereby improving both efficiency and solution quality. This collaborative paradigm, which embeds LLMs as components within EAs, defines the direction of low-level assistance in optimization. In the following, we systematically review research in this area, focusing on four key stages of EAs where LLMs have been applied: initialization, evolutionary operators, algorithm configuration, and evaluation.
@@ V-B1 Initialization
In the optimization process, the quality of the initial solution strongly influences both convergence efficiency and final performance. By leveraging their rich background knowledge, LLMs can provide high-quality initial solutions that embed prior information, thereby narrowing the search space and establishing a stronger foundation for subsequent iterations. Representative applications have demonstrated this potential across diverse domains. In Neural Architecture Search (NAS), Yu et al. [126] employed LLMs to propose plausible architectural components, introducing valuable prior knowledge before the search began. Jawahar et al. [128] used LLMs as predictors to guide the initialization phase of NAS more efficiently. Beyond NAS, Teukam et al. [132] applied LLM-based initialization in genetic algorithms for bioengineering, generating high-quality mutant libraries to accelerate enzyme design. In financial planning, De et al. [130] demonstrated that LLMs can generate viable initial portfolios for optimization tasks. Despite these advances, challenges remain. Zhao et al. [143] highlighted the high computational cost of using LLMs for initialization, as well as their difficulty in consistently satisfying strict constraints and generating feasible solutions, particularly in large-scale or highly complex problem settings.
@@ V-B2 Evolutionary Operators
Evolutionary operators are the core drivers of EAs. Before the advent of LLMs, researchers explored the use of neural networks, such as attention mechanisms and feed-forward networks [180], to simulate the behavior of individual operators or even entire algorithms. With the rise of iterative prompting approaches such as OPRO [28], attention has shifted toward employing LLMs directly to simulate operator behavior. This paradigm enables more adaptive search during optimization, as LLMs can exploit problem and solution characteristics without additional training, thereby surpassing the performance of conventional operators in some settings.
Several frameworks exemplify this direction. The LMX framework proposed by Meyerson et al. [125] employs few-shot prompting to use LLMs as intelligent operators for crossover and recombination of text-based genomes, generating semantically coherent variants without training. Liu et al. [127] extended this idea through the LMEA framework, where LLMs perform selection, crossover, and mutation under prompt guidance, coupled with a temperature self-adaptive mechanism to balance exploration and exploitation. Other studies have targeted specific operators or algorithms: PAIR [146] focuses on selection, LMPSO [148] adapts LLMs to PSO, and LEO [131] explores LLM-driven strategies for balancing exploration and exploitation.
This line of work has also expanded into multi-objective optimization. Liu et al. [129] were the first to empower MOEA/D [181] with zero-shot prompting, allowing LLMs to act as search operators and inspiring the design of more efficient white-box operators. To reduce interaction costs, Wang et al. [134] proposed a hybrid scheme in which LLMs generate only 10% of the population, while the rest are produced by traditional operators. Similarly, Liu et al. [141] called upon LLMs only when population improvement was insufficient, while relying on NSGA-II [182] otherwise. These hybrid approaches highlight the potential of selectively integrating LLMs to balance performance gains with computational efficiency.

### v3

@@ IV-C Challenges
With the growing application of LLMs in optimization, increasing efforts have been made to bridge natural language descriptions of optimization problems with formal mathematical models. Frameworks such as Chain-of-Experts [xiao2023chain] and ORLM [huang2025orlm] demonstrate that LLMs can not only complete modeling tasks efficiently but also uncover implicit constraints that may be overlooked by human experts, thus introducing a new paradigm for optimization modeling. Nevertheless, existing research still faces several key challenges:
  • Prompt-based methods are attractive for fast deployment and minimal data requirements, but the security story must distinguish public APIs from enterprise-grade isolated deployments. Public endpoints do pose leakage risks, whereas managed options (e.g., Azure OpenAI Service, Amazon Bedrock, Google Vertex AI) provide tenancy isolation so data remain within the customer boundary—crucial for healthcare and defense use. Even with isolation, current closed-source models still show gaps in formal mathematical reasoning and numerical rigor, so improving intrinsic reasoning and augmenting with validators remains a core technical challenge.
  • Learning-based methods such as ReSocratic [yang2024optibench] and OptMATH [lu2025optmath] pursue synthetic data and fine-tuning, but open-source by default is not universally warranted: closed-source models can also be fine-tuned on enterprise platforms with data isolation, and strong open models sometimes work zero-shot for certain modeling tasks. The harder issues are (i) scarcity and cost of high-quality, domain-specific datasets, (ii) repeat fine-tuning across domains/scales, (iii) long-term maintenance and regression control of tuned models, and (iv) the possibility that newer high-reasoning models may obviate some multi-stage or prompt-engineering pipelines. Advancing this line requires data-efficient synthesis, clearer criteria for when to fine-tune, and maintainable update protocols.
@@ V LLMs for Optimization Solving
Optimization solving is a core stage of the optimization workflow, and decades of research have produced a wide range of high-quality algorithms. In recent years, machine learning methods have been increasingly integrated into optimization to improve performance, and the advent of LLMs has further propelled this research into a new phase. In this section, we categorize existing work into three paradigms: LLMs as optimizers, low-level LLM-assisted optimization algorithms, and high-level LLM-assisted optimization algorithms. Section V-A introduces LLMs as optimizers, which directly employ LLMs to solve problems without relying on traditional frameworks (see Fig. 4). Sections V-B and V-C discuss how LLMs can enhance optimization algorithms at the component level and at the orchestration/design level, respectively (illustrated in Figs. 5 and 6). Section V-D then outlines the key challenges in this area. Representative studies are summarized in Table II in Supplementary Document.
@@ V-A LLMs as Optimizers
LLMs possess strong in-context learning capabilities, enabling them to solve optimization problems through iterative prompting. A pioneering work in this direction is the OPRO framework proposed by Yang et al. [yang2023large], which formulates optimization tasks in natural language and iteratively refines solutions based on historical trajectories. OPRO’s principal success lies in prompt-driven optimization: by exposing the model to solution–feedback trajectories, it leverages the LLM’s ability to infer patterns from context and generate semantically coherent candidate updates, which is particularly effective for discrete or language structured tasks and for rapid prototyping in human-in-the-loop workflows. Building on this paradigm, Guo et al. [guo2023towards] benchmarked LLMs on four classical optimization tasks i.e., gradient descent, hill climbing, grid search, and black-box optimization and proposed task-agnostic metrics to improve comparability. Beyond benchmarking, OPRO has been extended to multiple application domains. Jiang et al. [jiang2024large] applied it to combinatorial optimization for structural matrix ordering; Qiu et al. [qiu2024large] developed LMCO for wireless network design; Ghose et al. [ghose2025orfs] employed it for parameter tuning in open-source hardware workflows; and Jiang et al. [jiang2024optimizable] adapted the paradigm to jailbreak attack analysis, treating optimizable suffixes and harmfulness scores as part of the optimization trajectory.
@@
In contrast to OPRO’s direct use of solution–feedback trajectories, subsequent research has explored alternative forms of iteration to guide LLMs in optimization tasks. Lange et al. [lange2024large] sorted discretized candidate solutions from best to worst and provided the ordered sequence as context to the LLM. This reframes the task from pure solution generation to sequence recognition and reasoning, leveraging the model’s ability to capture ordering patterns. Cheng et al. [cheng2024trace] proposed the OPTO framework, which feeds LLMs with structured execution traces (e.g., intermediate variable values, function calls, and runtime outputs) to simulate backpropagation-like optimization and improve task understanding. Beyond textual iteration, multimodal extensions have further broadened the paradigm by incorporating visual information. Huang et al. [huang2025multimodal] addressed vehicle routing problems using both textual descriptions and map images to mimic human spatial cognition. Elhenawy et al. [elhenawy2024visual] solved the classic traveling salesman problem by combining scatter plot inputs with a multi-agent collaboration mechanism for path search. Zhao et al. [zhao2025bridging] extended this idea to graph-structured problems by converting abstract graphs into visual images, enabling multimodal LLMs to capture topological features in a manner more akin to human reasoning. Collectively, these studies demonstrate that extending OPRO with sequence-based, trace-based, or multimodal inputs can substantially enrich the optimization capabilities of LLMs.
Despite the progress demonstrated by iterative prompting frameworks, systematic empirical evaluations have revealed important performance boundaries. Zhang et al. [zhang2024revisiting] highlighted the strong model dependency of OPRO, showing that smaller models struggle to exploit long iterative histories effectively. More broadly, Huang et al. [huang2024exploring] systematically evaluated LLMs on both discrete and continuous black-box optimization tasks and confirmed their considerable strengths in symbolic and language-structured domains, while simultaneously exposing significant weaknesses in purely numerical settings such as continuous optimization. These findings suggest that the applicability of stand-alone LLM-based optimization is more constrained than initially anticipated. The performance gap widens considerably when problems lack natural language structure, when solutions demand precise numerical calibration, or when evaluation budgets are limited and sample efficiency becomes paramount.
@@
Based on these limitations, researchers have begun to explore alternative avenues for applying LLMs in optimization. One line of work attributes the drawbacks to prompt engineering and the inherent constraints of current models, and thus seeks to improve optimization capability through fine-tuning or training. Abgaryan et al. [abgaryan2024llms] fine-tuned an LLM to output solutions in a single step, bypassing multi-round iteration by training on instruction–solution pairs. Li et al. [li2024pretrained] introduced POM, a pretrained optimizer model that generalizes across black-box problems in a zero-shot manner. Chen et al. [chen2024generalists] proposed LLOME, a two-level optimization framework that incorporates preference learning to generate biologically valid sequences under complex biophysical constraints. Meanwhile, an increasing body of research suggests that LLMs may be more effective when used to assist rather than replace traditional optimization algorithms. This perspective motivates the next paradigm of low-level LLM-assisted optimization, where LLMs are embedded into existing frameworks to enhance their performance.
@@ V-B Low-level LLMs for Optimization Algorithms
There is a natural complementarity between the collective intelligence of EAs and the knowledge-driven reasoning of LLMs [chao2024large]. While EAs excel at heuristic exploration through population-based search, LLMs contribute powerful semantic understanding and generative capabilities that can guide this process, thereby improving both efficiency and solution quality. This collaborative paradigm, which embeds LLMs as components within EAs, defines the direction of low-level assistance in optimization. In the following, we systematically review research in this area, focusing on four key stages of EAs where LLMs have been applied: initialization, evolutionary operators, algorithm configuration, and evaluation.
@@ V-B1 Initialization
In the optimization process, the quality of the initial solution strongly influences both convergence efficiency and final performance. By leveraging their rich background knowledge, LLMs can provide high-quality initial solutions that embed prior information, thereby narrowing the search space and establishing a stronger foundation for subsequent iterations. Representative applications have demonstrated this potential across diverse domains. In Neural Architecture Search (NAS) [xue2021self, huang2025exploring], Yu et al. [yu2023gpt] employed LLMs to propose plausible architectural components, introducing valuable prior knowledge before the search began. Jawahar et al. [jawahar2023llm] used LLMs as predictors to guide the initialization phase of NAS more efficiently. Beyond NAS, Teukam et al. [teukam2024integrating] applied LLM-based initialization in genetic algorithms for bioengineering, generating high-quality mutant libraries to accelerate enzyme design. In financial planning, De et al. [de2023optimized] demonstrated that LLMs can generate viable initial portfolios for optimization tasks. Despite these advances, challenges remain. Zhao et al. [zhao2025can] highlighted the high computational cost of using LLMs for initialization, as well as their difficulty in consistently satisfying strict constraints and generating feasible solutions, particularly in large-scale or highly complex problem settings.
@@ V-B2 Evolutionary Operators
Evolutionary operators are the core drivers of EAs. Before the advent of LLMs, researchers explored the use of neural networks, such as attention mechanisms and feed-forward networks [lange2024evolution], to simulate the behavior of individual operators or even entire algorithms. With the rise of iterative prompting approaches such as OPRO [yang2023large], attention has shifted toward employing LLMs directly to simulate operator behavior. This paradigm enables more adaptive search during optimization, as LLMs can exploit problem and solution characteristics without additional training, thereby surpassing the performance of conventional operators in some settings.
Several frameworks exemplify this direction [huang2024advancing]. The LMX framework proposed by Meyerson et al. [meyerson2024language] employs few-shot prompting to use LLMs as intelligent operators for crossover and recombination of text-based genomes, generating semantically coherent variants without training. Liu et al. [liu2024large] extended this idea through the LMEA framework, where LLMs perform selection, crossover, and mutation under prompt guidance, coupled with a temperature self-adaptive mechanism to balance exploration and exploitation. Other studies have targeted specific operators or algorithms: PAIR [ali2025pair] focuses on selection, LMPSO [shinohara2025large] adapts LLMs to PSO, and LEO [brahmachary2025large] explores LLM-driven strategies for balancing exploration and exploitation.
This line of work has also expanded into multi-objective optimization. Liu et al. [liu2025large] were the first to empower MOEA/D [zhang2007moea] with zero-shot prompting, allowing LLMs to act as search operators and inspiring the design of more efficient white-box operators. To reduce interaction costs, Wang et al. [wang2024largeicic] proposed a hybrid scheme in which LLMs generate only 10% of the population, while the rest are produced by traditional operators. Similarly, Liu et al. [liu2024largemulti] called upon LLMs only when population improvement was insufficient, while relying on NSGA-II [deb2002fast] otherwise. These hybrid approaches highlight the potential of selectively integrating LLMs to balance performance gains with computational efficiency.